Two objects are attracted to each other by a gravitational force F. If one mass is doubled and the other is tripled without changing the distance, what is the new gravitational force between the objects in terms of F?

A. 4 F B. 9 F C. 6 F D. 1/9 F E. 1/6 F

Please explain answer in detail if possible.
Answer #1

gravitational force = G*m1*m2/ r2

If one mass is doubled and the other is tripled without changing the distance

m1=> 2*m1

and m2 = 3*m2

and r = r (same no change )

F ' = G*2*m1*3*m2 /r^2 = 6* G*m1*m2/ r2 = 6F

C. 6 F

Good Luck !! for exam
let me know in a comment if there is any problem or doubts please wait for me to respond.